# A Guide to Misinformation Detection Data and Evaluation

### v3

@@ Abstract.
journalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; August 03–07, 2025; Toronto, ON
@@ 3.3. Feasibility
This has 3 main implications. First, there is an urgent need for higher quality claims datasets. Second, supplementary information beyond claim text, such as claim dates, authors, and other additional context may help alleviate this problem. However, it is critical to assess how the specific information a predictive method processes impacts the feasibility of the data it is being trained and evaluated on, to confirm that performance margins between methods represent real progress and not progress in predicting noise. Thus, third, we propose that even in the more favorable retrieval-augmented setting, Evaluation Quality Assurance (Section 5) is paramount.
We finally turn to paragraph datasets. We find (Figure 6) that they have a higher feasibility rate than claims datasets. This result makes sense considering that having a greater volume of text allows for more context, which reduces the amount of ambiguity in classifying misinformation. Thus, while it can be more complex to assess veracity with longer inputs, evaluation in this setting may be paradoxically easier. Since recent advances in LLMs (and multimodal systems that can process other types of inputs) provide many options for working with longer data, prioritizing this and other context-rich settings may be beneficial for the field.
@@ B.3.1. Claims Feasibility Evaluation
We count a claim as AI-labeled “feasible” as long as the AI marked the claim as either feasible with search or as feasible- no search required.
@@ B.3.4. Human Annotator Team and Aggregation
Our annotating team included 8 human experts: 3 authors and 5 colleagues of the authors. Each annotator completed an approximately equal number of examples (2̃90).

### v4

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737437isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Natural language processingccs: Applied computing Law, social and behavioral sciences
@@ 3.3. Feasibility
We next turn to paragraph datasets. We find (Figure 6) that they have a higher feasibility rate than claims datasets. This result makes sense considering that having a greater volume of text allows for more context, which reduces the amount of ambiguity in classifying misinformation. Thus, while it can be more complex to assess veracity with longer inputs, evaluation in this setting may be paradoxically easier. Since recent advances in LLMs (and multimodal systems that can process other types of inputs) provide many options for working with longer data, prioritizing this and other context-rich settings may be beneficial for the field.
We also observe in Appendix B.4 that feasibility has a substantial relationship (e.g., 0.56 Pearson correlation) with the performance of a veracity prediction system across different datasets. We note a caveat that it appears very noisy within datasets, i.e., feasibility is not a strong predictor for whether an individual example’s veracity will be assessed correctly, This is likely because there are many reasons beyond feasibility that a prediction system can be correct or incorrect, such as data leakage, lucky search results or assumptions about context or predictions in general, etc.
Overall, there are 3 main implications. First, there is an urgent need for higher quality claims datasets. Second, supplementary information beyond claim text, such as claim dates, authors, extended paragraphs, or other additional context may help alleviate this problem. However, it is critical to assess how the specific information a predictive method processes impacts the feasibility of the data it is being trained and evaluated on, to confirm that performance margins between methods represent real progress and not progress in predicting noise. Thus, third, we propose that even in the more favorable retrieval-augmented setting, Evaluation Quality Assurance (Section 5) is paramount.
@@ 8. Conclusion
Acknowledgements
This project has been partially funded through the Canada CIFAR AI Chair program and the Centre for the Study of Democratic Citizenship. Kellin Pelrine was supported by funding from the Fonds de recherche du Quebéc.
Author Contributions
Camille Thibault led the survey and curation of all the datasets. Jacob-Junqi Tian led all LLM-based experiments. Both first authors contributed across diverse other areas of the project. Gabrielle Peloquin-Skulski made key contributions to dataset curation and unification, and the early direction of research. Taylor Lynn Curtis led the public website and repositories for the data from prototypes to completion, and contributed to writing. James Zhou led paragraph datasets. Florence Laflamme contributed human analysis and annotation. Luke Yuxiang Guan contributed to the website and data repository. Reihaneh Rabbany and Jean-François Godbout advised the project. Kellin Pelrine advised and managed the project. Pelrine also proposed foundational ideas, including each quality assessment, the flaw in current metrics, EQA, and the research direction overall.
@@ B.3.1. Claims Feasibility Evaluation
We count a claim as AI-labeled “feasible” as long as the AI marked the claim as either feasible with search or as feasible- no search required. We use GPT-4o-mini as the AI model.
@@ B.3.4. Human Annotator Team and Aggregation
Our annotating team included 8 human experts: 3 authors and 5 colleagues of the authors. Each annotator completed an approximately equal number of examples (2̃90). On the portion of the data labeled by multiple annotators (see below), the average pair-wise Cohen Kappa was 0.519 and percentage agreement 76.9%, indicating substantial agreement.
@@
B.4. Feasibility and Predictability
The veracity of infeasible examples, by definition, cannot be assessed definitively – but predictive systems can still make predictions for them, which may or may not turn out correct. In this section, we investigate empirically the connection between feasibility and correct predictions.
To do so, we sample 300 examples from the test set of every claims dataset, with the exception of six smaller datasets that have less than 300 samples in their test set, where we instead use the train set. For each of these examples, we perform veracity prediction using GPT-4o with OpenAI’s web search tool, with the following prompt (this is very similar to the other search-enabled predictive results that followed Tian et al. (2024), but a simpler and more recently available procedure for this more recent experiment):
Your task is to analyze the factuality of the given statement.
You may invoke the search tool as many times as needed to retrieve up-to-date information. However, before invoking the tool, you must explain your rationale for doing so.
After providing all your analysis steps, summarize your analysis and state True statement; Factuality: 1” if you think the statement is factual, or False statement; Factuality: 0” otherwise.
To match the GPT-4o predictive system used here, we likewise use GPT-4o to predict feasibility, following the same process described in Appendix B.3.1). GPT-4o predicts lower feasibility rates than GPT-4o-mini used in the main paper, but the relationships described below are similar for both models.
Figure 16. Performance (in macro F1 score) of feasibility at predicting whether veracity classification will be correct. Performance is limited for most datasets.
In Figure 16, we observe a distinct but limited relationship between feasibility and predictability at the example level. Averaged across the datasets, feasibility has 51.4 macro F1 in predicting whether the classifier will evaluate veracity correctly. Meanwhile, if we pool the data, feasibility then has 56.0% macro F1 in predicting veracity. This is above the 50% random threshold, so there is some correspondence. At the same time, it is clearly far from a standalone explanation of classifier success or failure.
Figure 17. Feasibility vs. Predictability, with trend line plotted by OLS with 95% confidence interval. There is a substantial positive correlation: more feasible datasets are more predictable.
However, if instead of trying to make assessments on individual examples, we compare the overall feasibility (i.e., proportion of feasible examples) and prediction correctness rates (measured by macro F1 of the predictions vs. ground truth veracity labels) of each dataset, then we observe a strong correspondence: there’s 0.563 Pearson correlation, and we observe the positive relationship in Figure 17.
Overall, it makes sense that the example level results are noisy: one can make a correct or incorrect prediction for myriad reasons, and even if it’s infeasible to reliably make a correct one, one can get lucky in interpreting it the right way, finding the right thing in the search results, data leakage, just rolling the dice successfully on a random prediction, etc. There’s also noise in the feasibility classifier, and even humans disagree around a quarter of the time on which examples are actually feasible (Appendix B.3.4).
Aggregate feasibility, meanwhile, gives quite a strong indicator of dataset difficulty. This can mean two things: (A) There is just a lot of noise at the example level, which is averaged out at the group (dataset) level, (B) both feasibility and performance reflect some other aspects of dataset difficulty and quality: e.g., complicated claims seem less feasible and are indeed harder to assess. We hypothesize based on these results that both these possibilities are at play, in other words, there is a causal within-group effect partially obscured by noise, plus confounding at group level. We caution that many complex and intersecting factors are at play between dataset characteristics and veracity prediction systems, so this result may not be representative in all scenarios. Nonetheless, we see how feasibility can be highly informative in assessing predictive performance and dataset difficulty.